log(sub5)(x+10)+log(sub5)(x-10)=3
solve for x.

Answer by stanbon(75887)

log(sub5)(x+10)+log(sub5)(x-10)=3
-------------------
log5[(x+10)(x-10)] = 3
----
x^2-100 = 5^3 = 125
-----
x^2 = 25
----
x = 5 or x = -5
----
Both are extraneous solutions
because both values would make
x-10 negative. And you can't
have the log of a negative.
----
Ans: No solution::
Cheers,
Stan H.